OpenStudy (anonymous):

Find the domain and range of f(x)=-2x+|3sinx|

OpenStudy (anonymous):

well i need your help to sole this question? will u?

OpenStudy (anonymous):

of course :)

OpenStudy (anonymous):

good.. now tell me is there any value of x which if u put in the f(x) will make it undefined?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

why? do u have a reason to validate your point?

OpenStudy (anonymous):

If you plug 0 in for x it makes the equation = 0

OpenStudy (anonymous):

what is the problem in that? 0 is a well known number & it is not undefined.. any else no. which can make the equation undefined i mean that is not in the realm of real numbers?

OpenStudy (anonymous):

Well ... our domain is \[(-\infty,\infty)\] you can think upon why..

OpenStudy (anonymous):

and the range is the same

OpenStudy (anonymous):

now we come to range.. that may smell a bit complicated but it is not at least in this question.. it is again \[(-\infty,\infty)\] you can imagine if u put a very small number in this equation u will get a large no. in return..& vice versa.. think upon it...